Solve for X.
ax+b=cx-d;2x+9=7x-1

Answer by swincher4391(1107)

ax + b = cx - d
Get common terms together.
ax - cx = -b - d
Factor out an x.
x(a-c) = -(b+d)
x = -(b+d) / (a-c) <---- answer
By our first answer, 2x + 9 = 7x -1 where a=2, b = 9 , c =7 and d = 1 then

-(9+1)/(2-7) = -10/-5 = 2
x = 2. <----- answer